# SymCERE: Symmetric Contrastive Learning for Robust Review-Enhanced Recommendation

### v2

@@ 3.3. Final Objective Function
The final training objective is a weighted sum of the cross-modal loss, the intra-modal loss, and a complementary Bayesian Personalized Ranking (BPR) loss for intra-modal regularization. The total loss is:
(13) $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cross-modal}}+\alpha\cdot\mathcal{L}_{\text{intra-modal}}+\beta\cdot\mathcal{L}_{\text{BPR}}$
where $\alpha$ and $\beta$ are hyperparameters that balance the contributions of the contrastive alignment tasks and the recommendation ranking task. In our experiments, we set $\alpha=0.5$ and $\beta=0.05$.
@@ 4. Experiments
We conducted a comprehensive set of experiments to evaluate the performance of our proposed model, SymCERE. The primary research questions we aim to answer are:
  • RQ1: Does our unified contrastive method outperform state-of-the-art GNN-based and self-supervised recommendation baselines?
  • RQ2: How does our approach to mitigating false negatives and popularity bias contribute to the model’s effectiveness? (Addressed via ablation studies in Section 6.1)
  • RQ3: What is the underlying semantic mechanism by which our model aligns graph and text modalities? (Addressed via a case study in Section 6.2)
@@ 4.1.1. Datasets
To ensure a robust and comprehensive evaluation, we curated a collection of 15 datasets from three distinct sources. The first two sources are standard e-commerce benchmarks: 13 categories from the Amazon Reviews dataset (McAuley et al., 2015; He and McAuley, 2016) and the Yelp Open Dataset (Yelp, 2024). These are suitable for their link between user-item interactions and English review texts (Zhao et al., 2023; Zhou et al., 2023). To assess the generalizability of our method beyond standard e-commerce benchmarks, we also used a semi-public travel site dataset. It contains public user reviews in Japanese (Rakuten Group, Inc., 2020) and private staying records. We extracted a subset filtered by user (posted at least two reviews) and the period corresponding to each posted review.
To ensure data quality, we followed a standard pre-processing protocol (He et al., 2017, 2020). For each dataset (except the semi-public travel site), we applied the 5-core setting, filtering out users and items with fewer than five interactions. This ensures that every user and item in the evaluation has a minimal history to learn from. The statistics of the processed datasets are summarized in Table 1.
@@ 4.1.2. Evaluation Protocol
For each user, we split their interaction history chronologically into training, validation, and test sets with a ratio of 80:10:10. This ensures that the model is evaluated on its ability to predict future interactions based on past behavior.
We adopt a top-K item ranking evaluation protocol. For each user in the test set, the model is tasked with ranking the ground-truth item against all other items in the catalogue that the user has not interacted with in the training set. This ”all-ranking” protocol provides a comprehensive and unbiased evaluation of the model’s performance, avoiding potential biases from negative sampling during evaluation (Krichene and Rendle, 2020; Rendle et al., 2020).
We evaluate model performance using two standard top-K ranking metrics (Järvelin and Kekäläinen, 2002; Herlocker et al., 2004):
  • HR@10 (Hit Ratio at 10): Measures the proportion of users for whom the ground-truth item is ranked within the top 10.
  • NDCG@10 (Normalized Discounted Cumulative Gain at 10): A position-aware metric that assigns higher scores to hits at higher ranks in the top-10 list.
4.1.3. Baseline Models
We compare our model against a suite of competitive baselines to establish a strong performance benchmark. These models represent a range of approaches, from classic neural models to state-of-the-art GNN and self-supervised methods:
  • NeuMF (He et al., 2017): A classic neural recommendation model combining matrix factorization and a multi-layer perceptron to capture both linear and non-linear user-item relationships.
  • SGL (Wu et al., 2021): A self-supervised learning approach that enhances GNNs by creating auxiliary training signals through graph augmentations and contrastive learning.
  • DGCLR (Li et al., 2022): A self-supervised graph framework that disentangles user and item representations into different semantic factors using a channel-aware contrastive learning objective.
  • NGCF (Wang et al., 2019): A GNN-based model that explicitly encodes high-order connectivity in the user-item interaction graph through embedding propagation.
  • LightGCN (He et al., 2020): A simplified GNN that has become a standard, strong baseline, removing non-linearities and feature transformations for efficiency and effectiveness in collaborative filtering.
@@ 4.1.4. Implementation Details
To ensure fair and reproducible comparisons, the baseline models, except for DGCLR, were implemented within the RecBole framework (Zhao et al., 2021). The proposed SymCERE model, its related models, and DGCLR are implemented in non-RecBole implementations, which are available anonymously on Github. For the text encoder, we used the pre-trained LLM ”sarashina2.2-0.5b-instruct” for the semi-public travel dataset and ”blair-roberta-base” for other datasets, using LoRA. All experiments were performed on an NVIDIA H100 GPU (90GB).
@@ 5.1. Overall Performance Comparison (RQ1)
5.1.1. Hit Ratio (HR@10) Performance
Across the HR@10 metric, our SymCERE method shows strong performance. Our model variants secure the top position in 13 out of the 15 datasets. For instance, on the Yelp dataset, our model with an NGCF backbone achieves an HR@10 of 0.3540, which represents a 34.4% relative improvement over the strongest baseline (DGCLR at 0.2560). Similarly, on the Pet Supplies dataset, our model achieves an HR@10 of 0.3860, a 38.7% relative improvement over the best-performing baseline, SGL. These results indicate that the integration of textual information via our contrastive method is effective at identifying relevant items for users.
5.1.2. Normalized Discounted Cumulative Gain (NDCG@10) Performance
A similar trend is observed for the NDCG@10 metric, which evaluates the ranking quality of the top-10 recommendations. Our multi-modal model with an NGCF backbone achieves the highest score in 13 out of 15 datasets. The improvements are substantial, highlighting the model’s ability to not only find relevant items but also rank them correctly. On the Yelp dataset, our model achieves an NDCG@10 of 0.112, a 43.6% relative improvement over the best baseline (NeuMF at 0.0788). On Cell Phones and Accessories, our model yields an NDCG@10 of 0.1780, a 18.4% relative improvement over the baseline NGCF. The performance gains are consistent across nearly all datasets, although the margin of improvement varies depending on the domain characteristics, as discussed in Section 6.
In summary, the empirical results support a positive answer to RQ1. Our proposed SymCERE method demonstrates a significant performance advantage over established collaborative filtering and self-supervised baselines across two standard ranking metrics. This indicates that our unified approach to mitigating false negatives and popularity bias can lead to more accurate and effective recommendations.
Table 2. Unified performance comparison using HR@10 and NDCG@10. This table contrasts standard baselines with our multi-modal method (SymCERE) applied to different GNN backbones. For each metric (row), the best result is highlighted in bold and the second-best is underlined.
Dataset Metric Standard Baselines Ours (SymCERE)
@@ 5.1.2. Normalized Discounted Cumulative Gain (NDCG@10) Performance
DGCLR NeuMF SGL NGCF LightGCN + NGCF + LightGCN
Clothing, Shoes & Jewelry HR@10 0.504 0.521 0.530 0.525 0.530 0.538 0.550 $\uparrow 3.8\%$
NDCG@10 0.357 0.415 0.412 0.420 0.411 0.425 0.398 $\uparrow 1.2\%$
CDs & Vinyl HR@10 0.335 0.453 0.456 0.450 0.419 0.488 0.495 $\uparrow 8.5\%$
NDCG@10 0.235 0.324 0.309 0.324 0.285 0.332 0.308 $\uparrow 2.6\%$
Tools & Home Improvement HR@10 0.248 0.330 0.336 0.323 0.334 0.402 0.412 $\uparrow 22.5\%$
NDCG@10 0.113 0.170 0.175 0.166 0.174 0.187 0.162 $\uparrow 6.7\%$
Sports & Outdoors HR@10 0.279 0.381 0.376 0.364 0.337 0.424 0.434 $\uparrow 14.0\%$
NDCG@10 0.135 0.209 0.206 0.202 0.191 0.226 0.195 $\uparrow 8.1\%$
Beauty & Personal Care HR@10 0.222 0.312 0.312 0.311 0.283 0.386 0.405 $\uparrow 29.8\%$
NDCG@10 0.085 0.141 0.135 0.138 0.124 0.162 0.135 $\uparrow 15.1\%$
Automotive HR@10 0.367 0.408 0.441 0.424 0.434 0.446 0.470 $\uparrow 6.6\%$
NDCG@10 0.210 0.260 0.281 0.278 0.276 0.285 0.265 $\uparrow 1.4\%$
Toys & Games HR@10 0.263 0.312 0.329 0.312 0.332 0.415 0.419 $\uparrow 26.2\%$
NDCG@10 0.125 0.170 0.174 0.175 0.173 0.198 0.173 $\uparrow 13.0\%$
Pet Supplies HR@10 0.192 0.268 0.279 0.268 0.234 0.373 0.387 $\uparrow 38.7\%$
NDCG@10 0.054 0.104 0.101 0.101 0.086 0.129 0.102 $\uparrow 24.2\%$
Health & Household HR@10 0.212 0.256 0.256 0.246 0.254 0.371 0.360 $\uparrow$ 44.9%
NDCG@10 0.064 0.106 0.108 0.102 0.106 0.132 0.107 $\uparrow$ 22.2%
Video Games HR@10 0.204 0.227 0.255 0.226 0.254 0.345 0.343 $\uparrow$ 35.3%
NDCG@10 0.074 0.099 0.109 0.098 0.104 0.132 0.109 $\uparrow$ 21.1%
Cell Phones & Accessories HR@10 0.289 0.254 0.306 0.296 0.299 0.411 0.411 $\uparrow 34.5\%$
NDCG@10 0.121 0.114 0.146 0.152 0.140 0.180 0.151 $\uparrow 18.3\%$
Patio, Lawn & Garden HR@10 0.340 0.216 0.358 0.351 0.354 0.433 0.450 $\uparrow 25.8\%$
NDCG@10 0.149 0.130 0.195 0.195 0.192 0.225 0.189 $\uparrow 15.4\%$
Office Products HR@10 0.316 0.195 0.351 0.347 0.351 0.430 0.430 $\uparrow 22.4\%$
NDCG@10 0.116 0.093 0.183 0.181 0.178 0.190 0.161 $\uparrow 3.9\%$
Yelp HR@10 0.256 0.238 0.193 0.182 0.146 0.344 0.285 $\uparrow$ 34.4%
NDCG@10 0.039 0.078 0.054 0.052 0.037 0.112 0.082 $\uparrow$ 43.6%
Travel HR@10 0.688 0.768 0.742 0.737 0.695 0.768 0.824 $\uparrow 7.3\%$
NDCG@10 0.266 0.268 0.296 0.269 0.279 0.269 0.380 $\uparrow 28.4\%$
Table 3. Ablation study for different model backbones. ‘w/o Norm‘ denotes the model without L2 normalisation, ‘$M$ w/ InfoNCE‘ is the model variant using InfoNCE for cross-modal loss, and ‘Max. Drop (%)‘ shows the performance degradation of the ‘$M$ w/ InfoNCE‘ model compared to the ‘Full‘ model.
Dataset Metric NGCF Backbone LightGCN Backbone
Full w/o Norm MM w/ InfoNCE Min. Drop (%) Full w/o Norm MM w/ InfoNCE Max. Drop (%)
Clothing, Shoes and Jewelry HR@10 0.538 0.002 0.536 $\downarrow 0.2\%$ 0.550 0.548 0.547 $\downarrow 0.6\%$
NDCG@10 0.425 0.001 0.428 $\uparrow 0.7\%$ 0.398 0.385 0.384 $\downarrow 3.6\%$
CDs and Vinyl HR@10 0.488 0.007 0.490 $\uparrow 0.4\%$ 0.495 0.492 0.463 $\downarrow 6.4\%$
NDCG@10 0.332 0.002 0.338 $\uparrow 1.8\%$ 0.308 0.303 0.295 $\downarrow 4.2\%$
Tools and Home Imp. HR@10 0.402 0.005 0.397 $\downarrow 1.3\%$ 0.412 0.409 0.396 $\downarrow 3.8\%$
NDCG@10 0.187 0.003 0.186 $\downarrow 0.3\%$ 0.162 0.156 0.153 $\downarrow 6.0\%$
Sports and Outdoors HR@10 0.424 0.002 0.426 $\uparrow 0.3\%$ 0.434 0.433 0.422 $\downarrow 2.9\%$
NDCG@10 0.226 0.001 0.227 $\uparrow 0.6\%$ 0.195 0.187 0.184 $\downarrow 5.6\%$
Beauty and Personal Care HR@10 0.386 0.001 0.390 $\uparrow 0.9\%$ 0.405 0.401 0.374 $\downarrow 7.7\%$
NDCG@10 0.162 0.000 0.158 $\downarrow 2.4\%$ 0.135 0.128 0.122 $\downarrow 9.6\%$
Automotive HR@10 0.446 0.005 0.450 $\uparrow 0.9\%$ 0.470 0.469 0.459 $\downarrow 2.3\%$
NDCG@10 0.285 0.002 0.283 $\downarrow 0.6\%$ 0.265 0.254 0.252 $\downarrow 4.9\%$
Toys and Games HR@10 0.415 0.001 0.417 $\uparrow 0.5\%$ 0.419 0.416 0.407 $\downarrow 3.0\%$
NDCG@10 0.198 0.000 0.197 $\downarrow 0.2\%$ 0.173 0.164 0.162 $\downarrow 6.3\%$
Pet Supplies HR@10 0.373 0.004 0.371 $\downarrow 0.7\%$ 0.387 0.383 0.369 $\downarrow 4.6\%$
NDCG@10 0.129 0.001 0.124 $\downarrow 3.9\%$ 0.102 0.096 0.093 $\downarrow 8.9\%$
Health and Household HR@10 0.358 0.002 0.355 $\downarrow$ 0.8% 0.357 0.343 0.291 $\downarrow$ 18.5%
NDCG@10 0.131 0.001 0.130 $\downarrow$ 0.8% 0.112 0.102 0.098 $\downarrow$ 12.5%
Video Games HR@10 0.332 0.006 0.331 $\downarrow$ 0.3% 0.326 0.280 0.222 $\downarrow$ 31.9%
NDCG@10 0.130 0.002 0.129 $\downarrow$ 0.8% 0.116 0.099 0.089 $\downarrow$ 23.3%
Cell Phones and Acc. HR@10 0.408 0.005 0.397 $\downarrow$ 2.7% 0.401 0.408 0.357 $\downarrow$ 11.0%
NDCG@10 0.178 0.001 0.170 $\downarrow$ 4.5% 0.161 0.146 0.149 $\downarrow$ 7.5%
Patio, Lawn and Garden HR@10 0.433 0.007 0.440 $\uparrow 1.6\%$ 0.450 0.440 0.447 $\downarrow 0.8\%$
NDCG@10 0.225 0.002 0.217 $\downarrow 3.7\%$ 0.189 0.179 0.182 $\downarrow 3.5\%$
Office Products HR@10 0.430 0.004 0.430 $\uparrow 0.0\%$ 0.430 0.430 0.422 $\downarrow 1.8\%$
NDCG@10 0.190 0.002 0.203 $\uparrow 6.9\%$ 0.161 0.147 0.147 $\downarrow 8.8\%$
Yelp HR@10 0.344 0.017 0.350 $\uparrow 1.7\%$ 0.293 0.224 0.210 $\downarrow 28.3\%$
NDCG@10 0.112 0.003 0.113 $\uparrow 0.9\%$ 0.089 0.070 0.068 $\downarrow 23.6\%$
Travel HR@10 0.768 0.040 0.789 $\uparrow 2.7\%$ 0.824 0.829 0.695 $\downarrow 15.7\%$
NDCG@10 0.269 0.006 0.298 $\uparrow 10.8\%$ 0.380 0.390 0.228 $\downarrow 40.0\%$

### v3

@@ 3.3. Final Objective Function
The final training objective is a weighted combination of these components:
(8) $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{cross-modal}}+\alpha\cdot\mathcal{L}_{\text{intra-modal}}+\beta\cdot\mathcal{L}_{\text{BPR}}+\lambda\|\Theta\|^{2}_{2}.$
where $\alpha=0.5$ and $\beta=0.05$ are hyperparameters determined empirically.
3.4. Theoretical Analysis: Geometric Mechanism of Semantic Anchoring
Our empirical results suggest that SymCERE exhibits “Semantic Anchoring,” prioritizing objective product attributes over subjective sentiment. We provide a geometric proof that this implies the model suppresses components that are non-discriminative with respect to the collaborative signal.
Figure 3. Geometric Mechanism of Semantic Anchoring. (a) Standard InfoNCE induces repulsion between semantically similar items (false negatives), causing diffusion on the hypersphere. This allows subjective noise ($\mathbf{v}_{\text{subj}}$) to persist as there is no structural pressure to align. (b) SymCERE removes these repulsive forces via false negative elimination, allowing items to form tight clusters around a common centroid $\boldsymbol{\mu}_{\mathcal{C}}$ aligned with objective features. (c) SemanticAnchoring: Under the strict L2 normalization constraint, maximizing alignment with the centroid forces the embedding to allocate its limited magnitude almost entirely to the objective component ($\lambda_{\text{obj}}\to 1$), effectively suppressing subjective noise ($\lambda_{\text{subj}}\to 0$).
3.4.1. Vector Decomposition on the Hypersphere
Consider a semantic embedding vector $\mathbf{t}$ on the unit hypersphere $\mathbb{S}^{d-1}$. We decompose $\mathbf{t}$ into a discriminative objective component ($\mathbf{v}_{\text{obj}}$) and a common-mode subjective component ($\mathbf{v}_{\text{subj}}$):
(9) $\mathbf{t}=\lambda_{\text{obj}}\mathbf{v}_{\text{obj}}+\lambda_{\text{subj}}\mathbf{v}_{\text{subj}},\quad\text{s.t.}\quad\|\mathbf{t}\|_{2}=1.$
3.4.2. Discriminative Orthogonality under Graph Constraints
Strictly removing false negatives via $\mathcal{N}_{i}$ forces the model to align the anchor with a set of semantically related items $\mathcal{C}$ (implicit cluster), maximizing the alignment with the cluster centroid $\boldsymbol{\mu}_{\mathcal{C}}$. Crucially, the collaborative graph links items based on specific functional substitutability (e.g., specific car parts), not generic sentiment. Thus, the centroid $\boldsymbol{\mu}_{\mathcal{C}}$ is defined by shared objective features:
(10) $\boldsymbol{\mu}_{\mathcal{C}}\propto\mathbf{v}_{\text{obj}}.$
Conversely, generic sentiment (e.g., ”good”) appears across diverse, unrelated clusters. While not necessarily zero-mean, it offers no contrastive power to distinguish cluster $\mathcal{C}$ from others. In the local tangent space of the optimal cluster, the subjective component is effectively orthogonal to the discriminative direction:
(11) $\mathbf{v}_{\text{subj}}\perp\boldsymbol{\mu}_{\mathcal{C}}.$
3.4.3. Signal Maximization under L2 Constraint
To maximize the inner product $\mathbf{t}^{\top}\boldsymbol{\mu}_{\mathcal{C}}$ under the constraint $\lambda_{\text{obj}}^{2}+\lambda_{\text{subj}}^{2}\approx 1$, the model must allocate magnitude solely to the component parallel to $\boldsymbol{\mu}_{\mathcal{C}}$. Allocating magnitude to $\mathbf{v}_{\text{subj}}$ wastes the limited L2 budget on a non-discriminative dimension:
(12) $\lim_{\text{training}\to\infty}\lambda_{\text{subj}}\to 0,\quad\lim_{\text{training}\to\infty}\lambda_{\text{obj}}\to 1.$
This proves that “Semantic Anchoring” is a structural consequence: L2 normalization forces the model to discard features—even strong ones like sentiment—that do not align with the fine-grained geometric structure imposed by the graph.
@@ 4. Experiments
We conducted comprehensive experiments to answer:
  • RQ1: Does SymCERE outperform state-of-the-art GNN-based and self-supervised recommendation baselines?
  • RQ2: How do false negative mitigation and geometric debiasing contribute to performance? (Ablation studies)
  • RQ3: What is the semantic mechanism driving the alignment? (Case study)
@@ 4.1.1. Datasets
We curated 15 datasets from three sources: 13 Amazon Reviews categories (McAuley et al., 2015), Yelp (Yelp, 2024), and a semi-public Travel dataset (Rakuten Group, Inc., 2020) to test generalizability across languages and domains. We applied 5-core filtering (except for Travel) to ensure data quality. Statistics are summarized in Table 1.
@@ 4.1.2. Evaluation Protocol
For each user, we split their interaction history chronologically into training, validation sets with a ratio of 80:20. This ensures that the model is evaluated on its ability to predict future interactions based on past behavior. We adopt a top-K item ranking evaluation protocol. For each user in the test set, the model is tasked with ranking the ground-truth item against all other items in the catalogue that the user has not interacted with in the training set. This ”all-ranking” protocol provides a comprehensive and unbiased evaluation. We evaluate model performance using two standard top-K ranking metrics (Järvelin and Kekäläinen, 2002; Herlocker et al., 2004): HR@10 (Hit Ratio at 10) and NDCG@10 (Normalized Discounted Cumulative Gain at 10).
4.1.3. Baselines and Comparison Scope
We compare our model against a suite of competitive baselines covering three categories. First, for collaborative filtering, NeuMF (He et al., 2017), NGCF (Wang et al., 2019), and LightGCN (He et al., 2020) represent the evolution from neural CF to graph-based methods. Second, regarding graph contrastive learning, SGL (Wu et al., 2021) serves as a baseline for self-supervised learning on user-item graphs. Third, in the category of multi-modal recommendation, we include SMORE (Ong and Khong, 2025) as a state-of-the-art multi-modal contrastive learning baseline. Comparing against SMORE specifically allows us to test the ”Fusion Gap” hypothesis—whether simple aggregation fails under noisy review data.
It is important to note that many recent Multi-Modal Recommendation (MMRec) works focus on Product Metadata (clean descriptions/images) (e.g., Amazon Product Data). In contrast, our work targets the more challenging and noisy domain of User Reviews (Amazon Review Data). Standard MMRec baselines designed for clean metadata are often ill-suited for the signal ambiguity inherent in reviews. Therefore, we compare against methods established in review-enhanced and collaborative filtering contexts to ensure a fair and rigorous evaluation of semantic alignment capabilities.
@@ 4.1.4. Implementation Details
To ensure fair and reproducible comparisons, the baseline models were implemented within the RecBole framework (Zhao et al., 2021). The proposed SymCERE model, its related models are implemented in non-RecBole implementations, which are available anonymously on Github. For the text encoder, we used the pre-trained LLM ”sarashina2.2-0.5b-instruct” for the semi-public travel dataset and ”blair-roberta-base” for other datasets, using LoRA. All experiments were performed on an NVIDIA H100 GPU (90GB).
Fairness in Multi-Modal Comparison.
To ensure a rigorously fair comparison, all multi-modal baselines (SMORE) utilise the exact same semantic embeddings generated by our text encoder from the user reviews. We deliberately do not use clean product metadata for baselines. This setup isolates the model architecture as the sole variable, testing each model’s capability to extract signals from noisy, unstructured user feedback (”Signal Ambiguity”).
SMORE Configuration.
To ensure fairness, SMORE utilized the same LLM backbones. Since SMORE requires item-level features, we adapted the data using an “Encode-Early” strategy: training reviews for item $v$ were individually encoded (CLS token) and aggregated via mean pooling. This process strictly adhered to the temporal constraints in Section 4.1.4 to prevent information leakage.
Temporal Leakage Prevention.
To prevent future information leaks, we enforce strict temporal constraints: For interaction $(u,v,t)$ at timestamp $t$, review text $r_{u,v}$ is the single review by $u$ for $v$ at $t$. We do not aggregate reviews or use reviews after $t$. All splits respect temporal ordering (80% earliest for training, 20% latest for testing).
@@ 5.1. Overall Performance Comparison (RQ1)
Table 2. Overall performance comparison (HR@10 and NDCG@10) against baselines. Best result is bolded; second-best is underlined. Max. Improv. (%) is calculated against the best performing baseline.
Dataset Metric GNN Baselines Multi-Modal Ours (SymCERE)
@@
NeuMF SGL NGCF LightGCN SMORE + NGCF + LightGCN
Clothing, Shoes & Jewelry HR@10 0.521 0.530 0.525 0.530 0.434 0.538 0.550 $\uparrow 3.8\%$
NDCG@10 0.415 0.412 0.420 0.411 0.311 0.425 0.398 $\uparrow 1.2\%$
CDs & Vinyl HR@10 0.453 0.456 0.450 0.419 0.322 0.488 0.495 $\uparrow 8.5\%$
NDCG@10 0.324 0.309 0.324 0.285 0.218 0.332 0.308 $\uparrow 2.6\%$
Tools & Home Improvement HR@10 0.330 0.336 0.323 0.334 0.181 0.402 0.412 $\uparrow 22.5\%$
NDCG@10 0.170 0.175 0.166 0.174 0.096 0.187 0.162 $\uparrow 6.7\%$
Sports & Outdoors HR@10 0.318 0.376 0.364 0.337 0.218 0.424 0.434 $\uparrow 14.0\%$
NDCG@10 0.209 0.206 0.202 0.191 0.121 0.226 0.195 $\uparrow 8.1\%$
Beauty & Personal Care HR@10 0.312 0.312 0.311 0.283 0.197 0.386 0.405 $\uparrow 29.8\%$
NDCG@10 0.141 0.135 0.138 0.124 0.078 0.162 0.135 $\uparrow 15.1\%$
Automotive HR@10 0.408 0.441 0.424 0.434 0.312 0.446 0.470 $\uparrow 6.6\%$
NDCG@10 0.260 0.281 0.278 0.276 0.195 0.285 0.265 $\uparrow 1.4\%$
Toys & Games HR@10 0.312 0.329 0.312 0.332 0.233 0.415 0.419 $\uparrow 26.2\%$
NDCG@10 0.170 0.174 0.175 0.173 0.115 0.198 0.173 $\uparrow 13.0\%$
Pet Supplies HR@10 0.268 0.279 0.268 0.234 0.174 0.373 0.387 $\uparrow 38.7\%$
NDCG@10 0.104 0.101 0.101 0.086 0.050 0.129 0.102 $\uparrow 24.2\%$
Health & Household HR@10 0.256 0.256 0.246 0.254 0.193 0.371 0.360 $\uparrow 44.9\%$
NDCG@10 0.106 0.108 0.102 0.106 0.063 0.132 0.107 $\uparrow 22.2\%$
Video Games HR@10 0.227 0.255 0.226 0.254 0.162 0.345 0.343 $\uparrow 35.3\%$
NDCG@10 0.099 0.109 0.098 0.104 0.063 0.132 0.109 $\uparrow 21.1\%$
Cell Phones & Accessories HR@10 0.254 0.306 0.296 0.299 0.296 0.411 0.411 $\uparrow 34.5\%$
NDCG@10 0.114 0.146 0.152 0.140 0.120 0.180 0.151 $\uparrow 18.3\%$
Patio, Lawn & Garden HR@10 0.216 0.358 0.351 0.354 0.372 0.433 0.450 $\uparrow 25.8\%$
NDCG@10 0.130 0.195 0.195 0.192 0.161 0.225 0.189 $\uparrow 15.4\%$
Office Products HR@10 0.195 0.351 0.347 0.351 0.355 0.430 0.430 $\uparrow 22.4\%$
NDCG@10 0.093 0.183 0.181 0.178 0.132 0.190 0.161 $\uparrow 3.9\%$
Yelp HR@10 0.238 0.193 0.182 0.146 0.074 0.344 0.285 $\uparrow 34.4\%$
NDCG@10 0.078 0.054 0.052 0.037 0.032 0.112 0.082 $\uparrow 43.6\%$
Travel (JP) HR@10 0.768 0.742 0.737 0.695 - 0.768 0.824 $\uparrow 7.3\%$
NDCG@10 0.268 0.296 0.269 0.279 - 0.269 0.380 $\uparrow 28.4\%$